# Kuramoto Orientation Diffusion Models

### v1

@@ 1 Introduction
Experiments on orientation-rich datasets, such as fingerprints, textures, and terrains, demonstrate that our Kuramoto orientation diffusion model consistently produces higher-fidelity images compared to standard diffusion baselines, often requiring fewer diffusion steps. Furthermore, our method achieves competitive results even on general image-generation benchmarks (e.g., CIFAR-10), highlighting its broader applicability beyond structured domains. Overall, this work bridges neural oscillation theory and modern score-based generative models, underscoring the potential of biologically inspired synchronization dynamics as structured priors, particularly beneficial for orientation-dense data.
@@ 3.1 Setup
Datasets, Baselines, and Metrics. We evaluate our proposed method across both general-purpose and orientation-rich image generation tasks. For standard benchmarking, we first test on CIFAR10 [34]. To specifically assess performance on orientation-dense data, we then apply our model to the SOCOFing fingerprint dataset [55], the Brodatz texture dataset [1, 8], and the ground terrain dataset [67]. The input image resolutions are $3{\times}32{\times}32$ for CIFAR10 and $1{\times}96{\times}96$ for SOCOFing. The Brodatz texture dataset originally consists of high-resolution samples depicting various textures (e.g., grass, water, sand, wool) with sizes of $1{\times}512{\times}512$ or $1{\times}1024{\times}1024$. Due to the limited number of available images, we increase the dataset size by dividing these high-resolution textures into smaller patches of $1{\times}32{\times}32$. Compared to Brodatz textures, the ground terrain dataset [67] contains a broader set of orientation-rich material textures (e.g., plastic, turf, steel, asphalt, leaves, stone, and brick) at the higher resolution of $128{\times}128$.
As a baseline, we primarily compare with the standard variance-preserving (VP) score-based generative model (SGM) [59]. To evaluate the quality of generated images, we use the Fréchet Inception Distance (FID) [21], a widely used metric for assessing visual fidelity and diversity. All models are trained until the FID score converges to ensure fair comparison.
@@ 5 Conclusion
Broader Impacts. Our approach may benefit socially relevant domains such as biometric security (e.g., synthetic fingerprint generation), medical imaging (e.g., fiber orientation modeling in MRI), and scientific visualization (e.g., materials analysis). While improved generative models can enhance capabilities in these areas, they also carry risks associated with the misuse of synthetic data. As with other generative technologies, we advocate for responsible research and deployment in accordance with ethical and societal guidelines.
@@ B.4 Navier-Stokes Fluid Velocity Field
Amplitude: $\frac{\mathop{d{\bm{z}}_{t}^{i}}}{\mathop{dt}}=-\frac{1}{2}\beta_{t}{\bm{z}}_{t}^{i}+\sqrt{\beta_{t}}\xi^{i},\quad{\bm{z}}_{0}=\log{{\bm{r}}_{0}}$ (21)
Phase: $\frac{d{\bm{\theta}}^{i}_{t}}{dt}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}K(t)\sin({\bm{\theta}}^{j}_{t}-{\bm{\theta}}^{i}_{t})+K_{\text{ref}}(t)\sin(\psi_{\text{ref}}-{\bm{\theta}}^{i}_{t})+\sqrt{2D_{t}}\xi^{i}$
@@
Following fluid–dynamics practice, we assess realism in the spectral domain at the final time step. From each 2D velocity field we compute the radial energy spectrum $E(k)$ and then: (i) fit a line to $\log E(k)$ vs. $\log k$ over the mid–wavenumber band $k\in[0.1,0.4]\cdot k_{max}$ (with $k_{max}=\min(H,W)/2$) and report the absolute slope difference, and (ii) compute the 1D Wasserstein distance between the normalized mean spectra of real and generated samples.
@@
SGM (VP-SDE) [59] 0.7435 0.0015
Ours (Kuramoto Phase + Log-VP Amplitude) 0.6954 0.0011
Table 9 shows that coupling phases through the local Kuramoto dynamics while diffusing the log–magnitude yields spectra closer to the true data: slope error is reduced by $6.5\%$ and Wasserstein distance is decreased by $27\%$ relative to a standard diffusion model that treats angle and magnitude independently. This improvement suggests that the non-isotropic and non-linear Kuramoto-based angular coupling can improve physical plausibility in vector–valued generative modeling, demonstrating the benefit of synchronization-based inductive biases in real angular data modeling.
@@ B.5 More Examples of Generative Samples
Figs. 1011,  12, and 13 present additional randomly generated samples from our 1000-step locally coupled Kuramoto orientation diffusion model on the SOCOFing fingerprint, Brodatz texture, ground terrain, and CIFAR10 datasets, respectively.

### v2

@@ 1 Introduction
Experiments on orientation-rich datasets, such as fingerprints, textures, and terrains, demonstrate that our Kuramoto orientation diffusion model consistently produces higher-fidelity images compared to standard diffusion baselines, often requiring fewer diffusion steps. Furthermore, our method remains competitive even on general CIFAR-10 benchmarks. Beyond images, we also report results on Earth/climate datasets on the 2D sphere and on Navier–Stokes fluid velocity fields, where the synchronization prior aligns with natural periodicity and angular structure, delivering consistent gains. Overall, this work bridges neural oscillation theory and modern score-based generative models, underscoring the potential of biologically inspired synchronization dynamics as structured priors.
@@ 3.1 Setup
Datasets, Baselines, and Metrics. We evaluate our proposed method across both general-purpose and orientation-rich image generation tasks. For standard benchmarking, we first test on CIFAR10 [35]. To specifically assess performance on orientation-dense data, we then apply our model to the SOCOFing fingerprint dataset [56], the Brodatz texture dataset [1, 8], and the ground terrain dataset [68]. The input image resolutions are $3{\times}32{\times}32$ for CIFAR10 and $1{\times}96{\times}96$ for SOCOFing. The Brodatz texture dataset originally consists of high-resolution samples depicting various textures (e.g., grass, water, sand, wool) with sizes of $1{\times}512{\times}512$ or $1{\times}1024{\times}1024$. Due to the limited number of available images, we increase the dataset size by dividing these high-resolution textures into smaller patches of $1{\times}32{\times}32$. Compared to Brodatz textures, the ground terrain dataset [68] contains a broader set of orientation-rich material textures (e.g., plastic, turf, steel, asphalt, leaves, stone, and brick) at the higher resolution of $128{\times}128$. We primarily compare with the standard variance-preserving (VP) score-based generative model (SGM) [60]. To evaluate the quality of generated images, we use the Fréchet Inception Distance (FID) [21], a widely used metric for assessing visual fidelity and diversity. All models are trained until the FID score converges to ensure fair comparison.
@@ 5 Conclusion
Broader Impacts. Our approach may benefit socially relevant domains such as biometric security (e.g., synthetic fingerprint generation), medical imaging (e.g., fiber orientation modeling in MRI), and scientific visualization (e.g., materials analysis). As with other generative technologies, we advocate for responsible research and deployment in accordance with ethical and societal guidelines.
Acknowledgments
We would like to thank anonymous reviewers for their constructive suggestions and feedback. Yue Song was supported by the gift donation from Cisco. T. Anderson Keller was supported by the Kempner Institute Research Fellowship. Takeru Miyato was supported by the ERC Starting Grant LEGO-3D (850533) and the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, project number: 276693517.
@@ B.4 Navier-Stokes Fluid Velocity Field
Amplitude: $\frac{\mathop{d{\bm{z}}_{t}^{i}}}{\mathop{dt}}=-\frac{1}{2}\beta_{t}{\bm{z}}_{t}^{i}+\boxed{\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}K(t)\cos({\bm{\theta}}^{j}_{t}-{\bm{\theta}}^{i}_{t})}+\sqrt{\beta_{t}}\xi^{i},\quad{\bm{z}}_{0}=\log{{\bm{r}}_{0}}$ (21)
Phase: $\frac{d{\bm{\theta}}^{i}_{t}}{dt}=\boxed{{\bm{r}}_{t}^{i}}\Big[\frac{1}{|\mathcal{N}_{i}|}\sum_{j\in\mathcal{N}_{i}}K(t)\sin({\bm{\theta}}^{j}_{t}-{\bm{\theta}}^{i}_{t})+K_{\text{ref}}(t)\sin(\psi_{\text{ref}}-{\bm{\theta}}^{i}_{t})\Big]+\sqrt{2D_{t}}\xi^{i}$
where the boxed terms couple the two channels: (i) local phase coherence accelerates amplitude growth (via $\cos$ in the $z$-SDE), and (ii) larger amplitude strengthens phase synchronization (amplitude factor ${\bm{r}}_{t}^{i}$ in the $\theta$-SDE). To avoid over-coupling late in diffusion (when noise dominates), we enable coupling only in the early stage, i.e., for $t<T/\alpha$ with a constant $\alpha>1$; for $t\geq T/\alpha$ the two SDEs evolve independently under their native log-VP/Kuramoto dynamics.
@@
Unconditional Generation. Following fluid–dynamics practice, we assess realism in the spectral domain at the final time step. From each 2D velocity field, we compute the radial energy spectrum $E(k)$ and then: (i) fit a line to $\log E(k)$ vs. $\log k$ over the mid–wavenumber band $k\in[0.1,0.4]\cdot k_{max}$ (with $k_{max}=\min(H,W)/2$) and report the absolute slope difference, and (ii) compute the 1D Wasserstein distance between the normalized mean spectra of real and generated samples.
@@
SGM (Amplitude-Phase Decomposition) 0.7435 0.0015
SGM (Cartesian Coordinates) 0.5590 0.0029
Ours (Naive Kuramoto Phase + Log-VP Amplitude) 0.6954 0.0011
Ours (Coupled Kuramoto Phase + Log-VP Amplitude) 0.3343 0.0005
Figure 10: The average energy spectra with fitted lines.
Table 9 shows that the Coupled Kuramoto Phase + Log-VP Amplitude model achieves the best spectral realism among all methods, with the lowest slope error and Wasserstein distance. Relative to the SGM (Amplitude–Phase Decomposition) baseline, this corresponds to a $55\%$ reduction in slope error and a $67\%$ reduction in Wasserstein distance. It also improves over SGM (Cartesian Coordinates) by $40\%$ in slope error and $83\%$ in Wasserstein, and over the Naive Kuramoto + Log-VP variant by $52\%$ (slope) and $55\%$ (Wasserstein). These gains indicate that explicitly coupling phase synchronization with amplitude evolution yields more physically plausible velocity spectra than treating angle and magnitude independently or modeling them without coupling. Fig. 10 plots the average energy spectra with fitted lines, where the coupled model shows a visibly tighter fit in the mid-frequency band and a closer overall spectral shape.
Conditional Forecasting. Building on the unconditional results above, we next consider a setting that is arguably even more useful in PDE modeling: forecasting future states from history. Unlike unconditional generation which tests realism in distribution, forecasting probes whether the model has learned the dynamics well enough to extrapolate in time.
We keep the coupled representations and train two conditioned score models ${\bm{s}}_{c}$ – one for phases $\bm{\theta}$} and one for log-magnitudes $\bm{z}$} – together with their unconditioned counterparts ${\bm{s}}_{u}$ via condition dropout. At inference, each pair is fused with classifier-free guidance (CFG) [22]:
Amplitude: $\tilde{{\bm{s}}}_{\omega}({\bm{z}})={\bm{s}}_{u}({\bm{z}})+\omega\big({\bm{s}}_{c}({\bm{z}}|c)-{\bm{s}}_{u}({\bm{z}})\big)$ (22)
Phase: $\tilde{{\bm{s}}}_{\omega}({\bm{\theta}})={\bm{s}}_{u}({\bm{\theta}})+\omega\big({\bm{s}}_{c}({\bm{\theta}}|c)-{\bm{s}}_{u}({\bm{\theta}})\big)$
where $c$ denotes the two-step Navier–Stokes history, and $\omega$ is a guidance temperature that controls adherence to the condition (larger $\omega$ denotes stronger conditioning). We evaluate accuracy using Mean Squared Error (MSE) between predicted and ground-truth velocity fields.
Table 10: MSE of Navier-Stokes fluid velocity predictions.
Methods MSE ($\downarrow$)
SGM (Cartesian Coordinates) 0.0260
Ours (Coupled Kuramoto Phase + Log-VP Amplitude) 0.0188
As shown in Table 10, the Coupled Kuramoto Phase + Log-VP Amplitude model reduces MSE from $0.0260$ to $0.0188$ compared with the SGM (Cartesian Coordinates) baseline, indicating more accurate conditional forecasts. Fig. 11 illustrates two examples; predictions closely match the ground truth in both horizontal and vertical velocity components.
Figure 11: Conditional one-step forecasts of our method on Navier–Stokes velocity: history (left), predictions (middle), and ground truth (right).
@@ B.5 More Examples of Generative Samples
Figs. 1213,  14, and 15 present additional randomly generated samples from our 1000-step locally coupled Kuramoto orientation diffusion model on the SOCOFing fingerprint, Brodatz texture, ground terrain, and CIFAR10 datasets, respectively.